# Spin structures and phase diagrams of the $\text{spin-}\frac{5}{2}$ triangular-lattice antiferromagnet Na2BaMn(PO4)2 under magnetic field

### v1

@@ III.3 Temperature and magnetic field phase diagrams
From the peak positions in the heat capacity data and the changes of the propagation vector we constructed the temperature and magnetic field phase diagrams shown in Figure 4. The phase boundaries for the spin polarized state for the two field directions are in agreement with a previous report [14]; however, our data suggest more complex HTH-T phase diagrams. For $\bm{H}\parallel\bm{\hat{c}}$, we evidence the formation of two pocket phases; a narrow between the AFM1 and the UUD phase, and AFM2. For $\bm{H}\perp\bm{\hat{c}}$ there is an absence of a phase transition for AFM2, a narrow region for $T<T_{N_{1}}$ where the AFM1 incommensurate ground state coexists with a commensurate phase, and a phase transition between a commensurate and incommensurate phase at higher magnetic fields.
The existence of a 1/3 plateau (UUD phase) is reported also in other classical TLA systems with $S=5/2$, e.g. Rb4Mn(MoO4)3 [38], Ba3MnNb2O9 [9] (Mn+2) and RbFe(MoO4)2 [22] (Fe+3). For classical systems the stability of the UUD phase in a finite temperature range indicates that thermal fluctuations are lifting the degeneracy of the ground state [39, 40]; otherwise, their absence would result in the appearance of the UUD phase at a single point. Thermal fluctuations are responsible for lowering the free energy of the system by selecting the highest entropic state, where two spins are aligned parallel and the other spin antiparallel to the direction of the external applied magnetic field. As the magnetic field further increases, a canted version of the UUD phase becomes stable (umbrella or V phase) [41, 42].
@@ III.4 Spin structures under magnetic field
The spin structure models of Na2BaMn(PO4)2 are based on the representation analysis of the modulation vector $\bm{k}=(1/3,1/3,k_{z})$, with the component $k_{z}$ varying throughout the phase diagram, as summarized in Section III.2. For all values of $k_{z}$, there are the same three irreducible representations of magnetic propagation vector group $\Gamma_{\mathrm{mag}}=\Gamma_{1}\oplus\Gamma_{2}\oplus\Gamma_{3}$, which are presented in Table 2. $\Gamma_{1}$ mode is a collinear arrangement of moments along the $c$-axis, with the up-up-down (UUD) amplitude pattern for zero phase, and plus-minus-zero for phase $\pi/4$. The modulation along the $c$-axis direction given by $k_{z}$ modulates the spin arrangement between these two patterns. $\Gamma_{2}$ has magnetic moments coplanar in the (001) plane, with moments forming the 120 structure along Mn–Mn bonds, and the $k_{z}$ modulation rotates the moments around the $c$-axis, showing right-handed helical modulation. $\Gamma_{3}$ is the same as $\Gamma_{2}$, but the $k_{z}$ modulation is left-handed.
Table 2: Irreducible representations of the magnetic moments of Mn ions in Na2BaMn(PO4)2. Irreps are derived for the Wyckoff position $1a$ in the space group $P\bar{3}$ and modulation vector $\bm{k}=(1/3,1/3,k_{z})$. Columns contain irrep symbols, normalized basis vectors and description of the spin structures.
irrep basis vector description
@@
Before presenting the results there are three important points to address regarding our measurements. First, unpolarized neutron diffraction can not distinguish between opposing chiralities of the magnetic structure, that is the difference between $\Gamma_{2}$ and $\Gamma_{3}$, but can determine the ratio of these mode amplitudes. Therefore, the magnetic structure can be described with reference to the in-plane component ($\Gamma_{2}$ and $\Gamma_{3}$ mixing) and out-of-plane component ($\Gamma_{1}$). Second, neutron diffraction cannot determine the global phase of the magnetic modulation, which is crucial to accurately depict the structure with a commensurate modulation vector. We have chosen the phase to best match the theoretical models presented in the next section. Finally, it is important to note that the refinement of the satellite reflection intensities allowed us to determine the antiferromagnetic component of the magnetic order. Under an applied magnetic field, there is an additional ferromagnetic component arising from magnetization that contributes to the nuclear reflections, however, determining its precise value from neutron diffraction measurements is unreliable. For depicting the magnetic structures under field we used the magnetization values determined in Ref. [14], also shown in Fig. 6(a). A comprehensive list of the refinement parameters, including mode amplitudes, ferromagnetic components, and global phases, is given in the Supplementary Material [43].
@@
Figure 5: Magnetic structures of Na2BaMn(PO4)2 based on single crystal neutron diffraction measurements under applied field. All structures are determined at 600 mK except e) AFM2 phase. Each structure is depicted from the directions along the $c$, $a$, and $a^{*}$-axis. Axes coordinates for these views are depicted only in a) showing $a_{1}$, $a_{2}$ and $c$-axis in red, green, blue respectively. a) At zero field and 600 mK, the spin configuration is incommensurate, co-planar with Y arrangement of spins and components along the $c$-axis and $[120]$ direction. e) At 1200 mK it is collinear along the $c$-axis, with moments in an up-up-down pattern. When a magnetic field is applied along the $c$-axis, b) spins gradually align with the field forming alternating YT layers at 1.2 T, c) UUD phase at 2.2 T, and d) umbrella V arrangement at 3.8 T. For the field applied along the $1\bar{1}0]$ direction the spins also gradually align along the field, f) with commensurately alternating V-W layers at 1.8 T, g) commensurately stacked W layers at 3 T, and h) incommensurately modulated V-W layers at 4.4 T. A detailed description of all phases is given in the main text.
@@
The spin arrangement at 600 mK and zero-field (AFM1 phase) shows in-plane and out-of-plane components, as displayed in Figure 5(a). The in-plane modes $\Gamma_{2}$ and $\Gamma_{3}$ sum up to a collinear component along the $[120]$ direction, which in combination with the out-of-plane mode $\Gamma_{1}$ results in a coplanar structure with Y arrangement of spin directions in the $z=0$ layer. The incommensurability of the modulation vector changes the phase of the spin configuration along the $[001]$ direction, resulting in cycloidal modulation of Mn spins with $\approx$60 interlayer phase shift and effective rotation of the Y configuration along the $c$-axis direction. The average magnitude of the individual moments is $3.70(8)\,\mu_{\mathrm{B}}$. Magnetic ordering at 1200 mK and zero field (AFM2 phase) is pure $\Gamma_{1}$ mode, that is, a collinear UUD structure along the $c$-axis, as shown in Figure 5(e). The absolute value of the magnetic moment is reduced to $\mu=2.07(1)\mu_{\mathrm{B}}$ and the structure resembles the 600 mK arrangement dynamically disordered within the $ab$-plane. Our refinements in zero-field match those reported in Ref. [23], where the authors determined the magnetic structures based on neutron powder diffraction of Na2BaMn(PO4)2 at 67 mK and 1.25 K, with small differences in the modulation vector.
@@
The field applied along the $c$-axis breaks the incommensurate modulation of the spin arrangement and introduces intermediate, commensurate phases before field-polarizing the spins [see Figures 5(b)-(d)]. At 1.2 T (YT phase) the refinement resembles the ground state Y configuration with spins slightly rotated along the field direction. The modulation along the $c$-axis alternates between the distorted Y configuration and T configuration, where the latter depicts two spins arranged antiferromagnetically within the $ab$-plane and one spin along the $c$-axis. Note, that the refinement at 1.2 T was poor compared to the other refinements (see Supplementary Material [43]), and should be treated as a rough approximation. Next, at 2.2 T spins arrange in the UUD configuration along the $c$-axis with a small in-plane component (UUD phase). The in-plane component is collinear with the $[120]$ direction, and five times smaller than the $c$-axis component. Finally, at 3.8 T the spins rearrange to an umbrella configuration (V phase), where all spins have positive components along the field, while the in-plane component spreads them around in a 120 configuration.
@@
With the field applied in-plane, all measured phases show almost coplanar arrangement, as the dominant component is $\Gamma_{1}$ mode with $c$-axis component, and uniform magnetization aligning the moments toward the $1\bar{1}0]$ direction. As a consequence, the spins are lying within the $(110)$ plane, that is, they have dominating components along the $c$-axis and field direction [see Figures 5(f)-(h)]. At 1.8 T the spins arrange in alternating V and W layers, with opening along the applied field direction $1\bar{1}0]$ (CVW phase). The V configuration is a superposition of UUD along the $c$-axis and uniform magnetization along the applied field resulting in spins out-of-plane, while the W configuration describes one spin along the field direction and the other two spread up and down with respect to the first one. When increasing the field to 3 T there are only W layers with commensurate modulation along the $c$-axis and opening of $\approx$120(CW phase). At 4.4 T the structure is incommensurate again and modulates between the W and V configuration (IVW phase) with opening of $\approx$85. The latter incommensurate configuration is also referred to as a fan phase [44].
@@ III.5 Spin model Hamiltonian
The localized magnetic moments in Na2BaMn(PO4)2 form a triangular lattice in the $ab$-plane, with a Mn–Mn stacking along the $c$-axis as described in Section III.4. The phase diagrams and the spin structure discussed previously indicate the dynamics of a frustrated antiferromagnetic triangular lattice [41, 14]. Initially, for Na2BaMn(PO4)2 we considered a 2D nearest-neighbor-only classical Heisenberg model with $J_{1}=-0.11$ meV and $S=5/2$. The significant single-ion anisotropy of Mn+2 ions points to an easy-axis anisotropy in the system. In order to reproduce the plateau in the reported magnetization measurements as a function of applied fields along the $c$-axis [14], $k^{c}=0.035$ meV was added to our model Hamiltonian (see Fig. 6). Our first attempt to reproduce the experimental results with this simple model is justified by the fact that two successive magnetic transitions at zero magnetic field are reported in other 2D triangular-lattice easy-axis AFM with high spin number [21, 45, 38].
Figure 6: Magnetization as a function of applied fields along and perpendicular to the easy-axis. Experimental data taken from Ref. [14]. Simulation (color circles) including single-ion anisotropy with easy-axis along $c$. The length of the plateau is proportional to the single-ion anisotropy.
@@
The Monte Carlo simulations with this 2D model resulted in the color maps of the specific heat as a function of temperature and applied field which are shown in Figs 7(a) and 7(b). These heat maps reveal the boundaries between the different phases of the model. At finite temperatures and in zero field, a temperature separation of $\sim$0.05 K is observed between the AFM1 and AFM2 phases, and the system becomes paramagnetic at $\sim$0.95 K. When the field is applied along the easy-axis $c$, we identify at least four regions of the phase-space diagram in Fig. 7(a). Through spin dynamics simulations, we determined the spin configuration in each phase at zero temperature. AFM1: The ground state at zero field corresponds to a noncollinear $120^{\circ}$ spin configuration where one of the three spins in the magnetic unit cell points along the easy-axis (out-of-plane). UUD: At 2 T, the system displays a collinear up-up-down spin configuration. V: This phase resembles the UUD state, although it is noncollinear, and it is marked by the rotations of the down spin towards the direction of the applied field. SP: At 6 T, the Zeeman energy dominates the exchange interactions and spin polarizes the system. In contrast, when the field is applied in-plane, only two regions (AFM1 and SP) are distinguishable through the specific heat color map, as shown in Fig. 7(b). Starting from the ground state at zero field, when the external field dominates the single-ion anisotropy, one spin starts to point parallel to the field. At 3 T, we observe that two spins are perpendicular to the field, antiparallel to each other, but aligned along the easy-axis, while the third spin remains parallel to the applied field. For higher fields, all spins continuously align with the field until the spin-polarized state is reached, slightly below 7 T. This phase sequence is consistent with the classical easy-axis triangular-lattice Heisenberg model [41].
@@
This simple 2D model already seems to capture some features and the overall shape of the experimentally determined phase diagrams. It is important to mention that the experimental phase diagrams shown in Figure 4 were obtained from a combination of neutron diffraction and specific heat measurements, while the theoretical phase diagrams were determined from the specific heat results from Monte Carlo simulations. Therefore, it is possible that the specific heat will not reveal all the field-induced transitions between the different phases. In addition, neutron diffraction data suggest that the magnetic order is characterized by an incommensurate wave-vector along the $c$-axis, as shown in Figs. 2 and 3, indicating that Na2BaMn(PO4)2 is a 3D easy-axis antiferromagnet with considerable interlayer couplings. This corresponds to a spin configuration based on the $120^{\circ}$ state, which offers a global rotation from one layer to the next. It is also worth mentioning that based on Fig. 3(b), at zero field intense magnetic Bragg peaks are observed at positions with magnetic propagation vectors $\bm{k}=(1/3,1/3,\pm k_{z})$. Propagation vectors of this form have been attributed to frustrated interlayer couplings and are reported in compounds with stacked magnetic triangular layers such as the isostructural RbFe(MoO4)2 [46].
@@
In order to include the interaction between layers in our model, we attempted to reproduce the spiraling spin configuration in Na2BaMn(PO4)2 via a frustration induced by two interlayer couplings. While the frustration 3D-coupling model was recently proposed for Na2BaCo(PO4)2 [37], our study is the first to apply it to Na2BaMn(PO4)2, demonstrating its relevance also for S = 5/2 systems. It is based on the competition between $J_{2}$, $J_{a}$, and $J_{b}$, see Fig. 1(d). According to this model, the propagation vector along $c$ is given by [37]:
@@
The resulting ground state spin configuration matches very well the experimentally determined one, with the exception that the model has a commensurable wavelength of 5 layers to allow for technical simplifications of the simulations. The specific-heat phase diagram for field out-of-plane is shown in Fig. 7(c), which is very similar to the 2D case, except for a new pocket created between the UUD and V phases. We note that at 2 T, with the 3D coupling, the spin configuration is similar to the up-up-down state. For a field in-plane, while the 2D model only featured a single ordered phase, the 3D coupling introduced several phase boundaries to the diagram, although some of them are very faint, as seen in Fig. 7 (d). It is worth mentioning that the 3D model results in $\sim$0.18 K. separation between the two successive zero-field magnetic transitions, and the paramagnetic state is observed at $\sim$1.18 K. The evolution of the spin configurations for the two applied field directions at zero temperature is also displayed in Figure 8, and a further comparison between the simulated and experimentally determined magnetic structures is presented in detail in the Supplementary Material [43]. At this stage a perfect agreement between the experimentally determined magnetic structures and the theoretically obtained ones cannot be achieved, however, several features are captured by the proposed 3D model.
Detailed inelastic neutron scattering studies on single crystals for both field directions and additional theoretical modeling would be crucial to determine the Hamiltonian of the Na2BaMn(PO4)2 system. Such measurements have been used to determine the microscopic Hamiltonian in the analogue RbFe(MoO4)2 [47], which also exhibits Y, UUD, and V phases and field-induced incommensurate stacking along the $c$-axis [48]. For example, an alternative scenario might involve an Ising-like XXZ Hamiltonian where (i) the width of the 1/3 UUD plateau in Fig. 4(a) will depend on the XXZ exchange anisotropy of the $J_{1}$ bond, and (ii) the separation between the two zero-field transitions reported at 1.13 and 1.28 K will rely on the XXZ nature and on the 3D couplings [49].

### v2

@@ III.3 Temperature and magnetic field phase diagrams
From the peak positions in the heat capacity data and the changes of the propagation vector we constructed the temperature and magnetic field phase diagrams shown in Figure 4. The phase boundaries for the spin polarized state for the two field directions are in agreement with a previous report [Kim_2022]; however, our data suggest more complex HTH-T phase diagrams. For $\bm{H}\parallel\bm{\hat{c}}$, we evidence the formation of two pocket phases; a narrow between the AFM1 and the UUD phase, and AFM2. For $\bm{H}\perp\bm{\hat{c}}$ there is an absence of a phase transition for AFM2, a narrow region for $T<T_{N_{1}}$ where the AFM1 incommensurate ground state coexists with a commensurate phase, and a phase transition between a commensurate and incommensurate phase at higher magnetic fields.
The existence of a 1/3 plateau (UUD phase) is reported also in other classical TLA systems with $S=5/2$, e.g. Rb4Mn(MoO4)3 [Ishii_2011], Ba3MnNb2O9 [Lee_2014] (Mn+2) and RbFe(MoO4)2 [Svistov_2006] (Fe+3). For classical systems the stability of the UUD phase in a finite temperature range indicates that thermal fluctuations are lifting the degeneracy of the ground state [Villain_1980, Henley1989]; otherwise, their absence would result in the appearance of the UUD phase at a single point. Thermal fluctuations are responsible for lowering the free energy of the system by selecting the highest entropic state, where two spins are aligned parallel and the other spin antiparallel to the direction of the external applied magnetic field. As the magnetic field further increases, a canted version of the UUD phase becomes stable (umbrella or V phase) [Seabra_2011, Yamamoto_2014].
@@ III.4 Spin structures under magnetic field
The spin structure models of Na2BaMn(PO4)2 are based on the representation analysis of the modulation vector $\bm{k}=(1/3,1/3,k_{z})$, with the component $k_{z}$ varying throughout the phase diagram, as summarized in Section III.2. For all values of $k_{z}$, there are the same three irreducible representations (irreps) of the magnetic propagation vector group $\Gamma_{\mathrm{mag}}=\Gamma_{1}\oplus\Gamma_{2}\oplus\Gamma_{3}$, which are presented in Table 2. $\Gamma_{1}$ mode is a collinear arrangement of moments along the $c$-axis, with the up-up-down (UUD) amplitude pattern for zero phase, and plus-minus-zero for phase $\pi/4$. The modulation along the $c$-axis direction given by $k_{z}$ modulates the spin arrangement between these two patterns. $\Gamma_{2}$ has magnetic moments coplanar in the (001) plane, with moments forming the 120 structure along Mn–Mn bonds, and the $k_{z}$ modulation rotates the moments around the $c$-axis, showing right-handed helical modulation. $\Gamma_{3}$ is the same as $\Gamma_{2}$, but the $k_{z}$ modulation is left-handed. An explicit mode decomposition and notation are given in the Supplementary Material (see Eq. S1 [Supplemental]).
Table 2: Irreducible representations of the magnetic moments of Mn ions in Na2BaMn(PO4)2. Irreps are derived for the Wyckoff position $1a$ in the space group $P\bar{3}$ and modulation vector $\bm{k}=(1/3,1/3,k_{z})$. Columns contain irreps symbols, normalized basis vectors and description of the spin structures.
irreps basis vector description
@@
Before presenting the results there are three important points to address regarding our measurements. First, unpolarized neutron diffraction can not distinguish between opposing chiralities of the magnetic structure, that is the difference between $\Gamma_{2}$ and $\Gamma_{3}$, but can determine the ratio of these mode amplitudes. Therefore, the magnetic structure can be described with reference to the in-plane component ($\Gamma_{2}$ and $\Gamma_{3}$ mixing) and out-of-plane component ($\Gamma_{1}$). Second, neutron diffraction cannot determine the global phase of the magnetic modulation, which is crucial to accurately depict the structure with a commensurate modulation vector. We have chosen the phase to best match the theoretical models presented in the next section. Finally, it is important to note that the refinement of the satellite reflection intensities allowed us to determine the antiferromagnetic component of the magnetic order. Under an applied magnetic field, there is an additional ferromagnetic component arising from magnetization that contributes to the nuclear reflections, however, determining its precise value from neutron diffraction measurements is unreliable. For depicting the magnetic structures under field we used the magnetization values determined in Ref. [Kim_2022], also shown in Fig. 6. A comprehensive list of the refinement parameters, including mode amplitudes, ferromagnetic components, and global phases, is given in the Supplementary Material (see Table SII, and goodness-of-fit plots in Fig. S2 [Supplemental]).
@@
Figure 5: Magnetic structures of Na2BaMn(PO4)2 based on single crystal neutron diffraction measurements under applied field. All structures are determined at 600 mK except the AFM2 phase (e). Each structure is depicted from the directions along the $c$, $a$, and $a^{*}$-axis. Axes coordinates for these views are depicted only in (a) showing $a_{1}$, $a_{2}$ and $c$-axis in red, green, blue respectively. (a) At zero field and 600 mK, the spin configuration is incommensurate, co-planar with Y arrangement of spins and components along the $c$-axis and $[120]$ direction. (e) At 1200 mK it is collinear along the $c$-axis, with moments in an up-up-down pattern. When a magnetic field is applied along the $c$-axis, spins gradually align with the field forming (b) alternating YT layers at 1.2 T, (c) the UUD phase at 2.2 T, and (d) the umbrella V arrangement at 3.8 T. For the field applied along the $1\bar{1}0]$ direction the spins also gradually align along the field, with (f) commensurately alternating V-W layers at 1.8 T, (g) commensurately stacked W layers at 3 T, and (h) incommensurately modulated V-W layers at 4.4 T. A detailed description of all phases is given in the main text; numerical values used to render all panels are listed in the Supplementary Material in Table SII [Supplemental].
@@
The spin arrangement at 600 mK and zero-field (AFM1 phase) shows in-plane and out-of-plane components, as displayed in Figure 5(a). The in-plane modes $\Gamma_{2}$ and $\Gamma_{3}$ sum up to a collinear component along the $[120]$ direction, which in combination with the out-of-plane mode $\Gamma_{1}$ results in a coplanar structure with Y arrangement of spin directions in the $z=0$ layer. The incommensurability of the modulation vector changes the phase of the spin configuration along the $[001]$ direction, resulting in cycloidal modulation of Mn spins with $\approx$60 interlayer phase shift and effective rotation of the Y configuration along the $c$-axis direction. The average magnitude of the individual moments is $3.70(8)\,\mu_{\mathrm{B}}$. Magnetic ordering at 1200 mK and zero field (AFM2 phase) is pure $\Gamma_{1}$ mode, that is, a collinear UUD structure along the $c$-axis, as shown in Figure 5(e). The absolute value of the magnetic moment is reduced to $\mu=2.07(1)\mu_{\mathrm{B}}$ and the structure resembles the 600 mK arrangement dynamically disordered within the $ab$-plane. Our refinements in zero-field match those reported in Ref. [Zhang_2024], where the authors determined the magnetic structures based on neutron powder diffraction of Na2BaMn(PO4)2 at 67 mK and 1.25 K, with small differences in the modulation vector.
@@
The field applied along the $c$-axis breaks the incommensurate modulation of the spin arrangement and introduces intermediate, commensurate phases before field-polarizing the spins [see Figures 5(b)-(d)]. At 1.2 T (YT phase) the refinement resembles the ground state Y configuration with spins slightly rotated along the field direction. The modulation along the $c$-axis alternates between the distorted Y configuration and T configuration, where the latter depicts two spins arranged antiferromagnetically within the $ab$-plane and one spin along the $c$-axis. Note, that the refinement at 1.2 T was poor compared to the other refinements (see Fig. S2(b) in Supplementary Material [Supplemental]), and should be treated as a rough approximation. Next, at 2.2 T spins arrange in the UUD configuration along the $c$-axis with a small in-plane component (UUD phase). The in-plane component is collinear with the $[120]$ direction, and five times smaller than the $c$-axis component. Finally, at 3.8 T the spins rearrange to an umbrella configuration (V phase), where all spins have positive components along the field, while the in-plane component spreads them around in a 120 configuration.
@@
With the field applied in-plane, all measured phases show almost coplanar arrangement, as the dominant component is $\Gamma_{1}$ mode with $c$-axis component, and uniform magnetization aligning the moments toward the $1\bar{1}0]$ direction. As a consequence, the spins are lying within the $(110)$ plane, that is, they have dominating components along the $c$-axis and field direction [see Figures 5(f)-(h)]. At 1.8 T the spins arrange in alternating V and W layers, with opening along the applied field direction $1\bar{1}0]$ (CVW phase). The V configuration is a superposition of UUD along the $c$-axis and uniform magnetization along the applied field resulting in spins out-of-plane, while the W configuration describes one spin along the field direction and the other two spread up and down with respect to the first one. When increasing the field to 3 T there are only W layers with commensurate modulation along the $c$-axis and opening of $\approx$120(CW phase). At 4.4 T the structure is incommensurate again and modulates between the W and V configuration (IVW phase) with opening of $\approx$85. The latter incommensurate configuration is also referred to as a fan phase [Nagamiya_1968].
@@ III.5 Spin model Hamiltonian
The localized magnetic moments in Na2BaMn(PO4)2 form a triangular lattice in the $ab$-plane, with a Mn–Mn stacking along the $c$-axis as described in Section III.4. The phase diagrams and the spin structure discussed previously indicate the dynamics of a frustrated antiferromagnetic triangular lattice [Seabra_2011, Kim_2022]. Initially, for Na2BaMn(PO4)2 we considered a two-dimensional (2D) nearest-neighbor-only classical Heisenberg model with $J_{1}=-0.11$ meV and $S=5/2$. The significant single-ion anisotropy of Mn+2 ions points to an easy-axis anisotropy in the system. In order to reproduce the plateau in the reported magnetization measurements as a function of applied fields along the $c$-axis [Kim_2022], $k^{c}=0.035$ meV was added to our model Hamiltonian (see Fig. 6). Our first attempt to reproduce the experimental results with this simple model is justified by the fact that two successive magnetic transitions at zero magnetic field are reported in other 2D triangular-lattice easy-axis AFM with high spin number [Kitazawa_1999, Kimura_2008, Ishii_2011].
Figure 6: Magnetization as a function of applied fields along and perpendicular to the easy-axis. Experimental data based on Ref. [Kim_2022]. Simulation (color circles) including single-ion anisotropy with easy-axis along $c$. The length of the plateau is proportional to the single-ion anisotropy.
@@
The Monte Carlo simulations with this 2D model resulted in the color maps of the specific heat as a function of temperature and applied field which are shown in Figs 7(a) and 7(b). These heat maps reveal the boundaries between the different phases of the model. At finite temperatures and in zero field, a temperature separation of $\sim$0.05 K is observed between the AFM1 and AFM2 phases, and the system becomes paramagnetic at $\sim$0.95 K. When the field is applied along the easy-axis $c$, we identify at least four regions of the phase-space diagram in Fig. 7(a). Through spin dynamics simulations, we determined the spin configuration in each phase at zero temperature. AFM1: The ground state at zero field corresponds to a noncollinear $120^{\circ}$ spin configuration where one of the three spins in the magnetic unit cell points along the easy-axis (out-of-plane). UUD: At 2 T, the system displays a collinear up-up-down spin configuration. V: This phase resembles the UUD state, although it is noncollinear, and it is marked by the rotations of the down spin towards the direction of the applied field. SP: At 6 T, the Zeeman energy dominates the exchange interactions and spin polarizes the system. In contrast, when the field is applied in-plane, only two regions (AFM1 and SP) are distinguishable through the specific heat color map, as shown in Fig. 7(b). Starting from the ground state at zero field, when the external field dominates the single-ion anisotropy, one spin starts to point parallel to the field. At 3 T, we observe that two spins are perpendicular to the field, antiparallel to each other, but aligned along the easy-axis, while the third spin remains parallel to the applied field. For higher fields, all spins continuously align with the field until the spin-polarized state is reached, slightly below 7 T. This phase sequence is consistent with the classical easy-axis triangular-lattice Heisenberg model [Seabra_2011].
@@
This simple 2D model already seems to capture some features and the overall shape of the experimentally determined phase diagrams. It is important to mention that the experimental phase diagrams shown in Figure 4 were obtained from a combination of neutron diffraction and specific heat measurements, while the theoretical phase diagrams were determined from the specific heat results from Monte Carlo simulations. Therefore, it is possible that the specific heat will not reveal all the field-induced transitions between the different phases. In addition, neutron diffraction data suggest that the magnetic order is characterized by an incommensurate wave-vector along the $c$-axis, as shown in Figs. 2 and 3, indicating that Na2BaMn(PO4)2 is a three-dimensional (3D) easy-axis antiferromagnet with considerable interlayer couplings. This corresponds to a spin configuration based on the $120^{\circ}$ state, which offers a global rotation from one layer to the next. It is also worth mentioning that based on Fig. 3(b), at zero field intense magnetic Bragg peaks are observed at positions with magnetic propagation vectors $\bm{k}=(1/3,1/3,\pm k_{z})$. Propagation vectors of this form have been attributed to frustrated interlayer couplings and are reported in compounds with stacked magnetic triangular layers such as the isostructural RbFe(MoO4)2 [Hearmon_2012].
@@
In order to include the interaction between layers in our model, we attempted to reproduce the spiraling spin configuration in Na2BaMn(PO4)2 via a frustration induced by two interlayer couplings. While the frustration 3D-coupling model was recently proposed for Na2BaCo(PO4)2 [Woodland_2025], our study is the first to apply it to Na2BaMn(PO4)2, demonstrating its relevance also for S = 5/2 systems. It is based on the competition between $J_{2}$, $J_{a}$, and $J_{b}$, see Fig. 1(d). According to this model, the propagation vector along $c$ is given by [Woodland_2025]:
@@
The resulting ground state spin configuration matches very well the experimentally determined one, with the exception that the model has a commensurable wavelength of 5 layers to allow for technical simplifications of the simulations. The specific-heat phase diagram for field out-of-plane is shown in Fig. 7(c), which is very similar to the 2D case, except for a new pocket created between the UUD and V phases. We note that at 2 T, with the 3D coupling, the spin configuration is similar to the up-up-down state. For a field in-plane, while the 2D model only featured a single ordered phase, the 3D coupling introduced several phase boundaries to the diagram, although some of them are very faint, as seen in Fig. 7 (d). It is worth mentioning that the 3D model results in $\sim$0.18 K. separation between the two successive zero-field magnetic transitions, and the paramagnetic state is observed at $\sim$1.18 K. The evolution of the spin configurations for the two applied field directions at zero temperature is also displayed in Figure 8, and a further comparison between the simulated and experimentally determined magnetic structures is presented in detail in the Supplementary Material (see Fig. S3 [Supplemental]). At this stage a perfect agreement between the experimentally determined magnetic structures and the theoretically obtained ones cannot be achieved, however, several features are captured by the proposed 3D model.
Detailed inelastic neutron scattering studies on single crystals for both field directions and additional theoretical modeling would be crucial to determine the Hamiltonian of the Na2BaMn(PO4)2 system. Such measurements have been used to determine the microscopic Hamiltonian in the analogue RbFe(MoO4)2 [White_2013], which also exhibits Y, UUD, and V phases and field-induced incommensurate stacking along the $c$-axis [Sakhratov_2022]. For example, an alternative scenario might involve an Ising-like XXZ Hamiltonian where (i) the width of the 1/3 UUD plateau in Fig. 4(a) will depend on the XXZ exchange anisotropy of the $J_{1}$ bond, and (ii) the separation between the two zero-field transitions reported at 1.13 and 1.28 K will rely on the XXZ nature and on the 3D couplings [Jia_2024].
Our reported HTH-T phase diagrams in Na2BaMn(PO4)2 largely follow the archetypal frustrated triangular-lattice antiferromagnet scenario (coplanar Y and V phases separated by the UUD state) as summarized in Ref. [Starykh_2015]. A key difference from the strictly 2D picture is that the propagation vector acquires a field-tunable out-of-plane component, $k_{z}(H,T)\neq 0$, which signals weak and frustrated interlayer exchange and is captured by our minimal 3D model. We also note that the very low-field region of the TLA phase diagram is explicitly flagged as unsettled in Ref. [Starykh_2015] and is where we capture the AFM2 phase. Its character, i.e., a longitudinal (easy-axis) ordering that precedes the coplanar AFM1 on cooling, is consistent with the XXZ triangular model in the presence of weak easy-axis anisotropy and weak 3D coupling, and with theoretical expectations for zero field [Kawamura_2010]. Within this context, RbFe(MoO4)2 realizes the canonical Y–UUD–V sequence with commensurate stacking for field applied in-plane [Sakhratov_2022], whereas in CuCrO2 weak interlayer couplings and easy-axis anisotropy drive incommensurate stacking and a rich field-induced cascade [Mun_2014, Sakhratov_2016]; see also Ref. [Frontzek_2011] for the crucial role of interlayer exchange in the incommensurate phases. In Na2BaMn(PO4)2, the combination of easy-axis anisotropy and frustrated interlayer exchange could explain both the evolution $k_{z}\neq 0$ in applied field and the finite zero-field AFM2 pocket.